give the y values for x=-2,-1,0,1,2,3 for the function y=1/x-1
Found 2 solutions by pradhyumna.kumar@gmail.com, rmromero:
Answer by pradhyumna.kumar@gmail.com(62)

give the y values for x=-2,-1,0,1,2,3 for the function y=1/x-1
sol :: the values of y are - 0.333 ,- 0.5 , -1 , 0 , 1 ,0.5 .

Answer by rmromero(383)

 
Give the y values for x=-2,-1,0,1,2,3 for the function y=1/x-1
Answer:
x -2 -1 0 1 2 3
y -1/3 -1/2 -1 undefined 1 1/2

Show work in this space.

1
y = ______
x -1
When x = -2 When x = -1 When x = 0
1 1 1
y = ______ y = ______ y = ______
-2 - 1 -1 - 1 0 - 1

1 1 1
= ___ = _____ = _____ or -1
-3 -2 -1

When x = 1 When x = 2 When x = 3
1 1 1
y = _____ y = _______ y = ______
1 - 1 2 - 1 3 - 1

1 1 1
= ___ or undefined = ___ or 1 = _____
0 1 2